OpenStudy (anonymous):

Which method would be the simplest way to solve the system? 7x+5y=19 -7x-2y=-16 a. graphing b. substitution c. elimination d. distributive

OpenStudy (a1234):

Elimination, because you have -7 and 7, so you can just eliminate one of those equations.

OpenStudy (anonymous):

c. Elimination cuz u can eliminate -7 & 7

OpenStudy (anonymous):

What is the solution to the system 7x+5y=19 -7x-2y=-16 a.(2,1) b.(1,2) c.(7,5) d.(-7,-2)

OpenStudy (anonymous):

Wait let me work out then I'll tell u

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Its b.(1,2) cuz 7x+5y=19 -7x-2y=-16 -7x and 7x gets eliminated Then +5y-2y=3y 19-16=3 So 3y=3 Y=3/3 Y=1 So subtitute y=1 in to 7x +5y=19 7x +5(1)=19 7x=19-5 7x=2 X=14/7 X=2

OpenStudy (anonymous):

so b?

OpenStudy (anonymous):

Sorry 7x=14 (not 2)

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Its b

OpenStudy (anonymous):

can u help me with 4 more?

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

which method would be most efficient to solve the system? y=1/2x 2x+3y=28 a. graphing b. substitution c. elimination d. distributive

OpenStudy (anonymous):

Y=1/2x (is it 1 divide by 2x?)

OpenStudy (anonymous):

I Do not understand the question